OpenStudy (disneyvhs):

Triangle ABC has vertices A(-4,4), B(6,0), C(-4,0). Is triangle ABC a right triangle?

OpenStudy (danjs):



OpenStudy (danjs):

YES

OpenStudy (disneyvhs):



OpenStudy (mathmath333):

https://www.desmos.com/calculator/xcbodd3cq6 u can draw the lines

OpenStudy (anonymous):

first we find the length of all the AB BC AC; \[ds = \sqrt{((x1-x2)^2 +(y1-y2)^2)}\]

OpenStudy (anonymous):

the length of AB if we take A(-4,4) B(6,0) :- ds = sqrt(6+4)^2 + (0-4)^2) = sqrt(116)

OpenStudy (anonymous):

the length of BC , B(6,0) C(-4,0) ds = sqrt((6+4)^2 -0) = 10

OpenStudy (anonymous):

the length of AC , C(-4,0) A(-4,4); ds =sqrt(( -4 - -4)^2 + (4-0)^2) = 4;

OpenStudy (anonymous):

you have AB = sqrt(116) AC = 4 and BC = 10; to proof that this triangle you need to proof that every two of edges bigger than the third

OpenStudy (anonymous):

AB + BC > AC sqrt(116) +10 > 4t his correct; BC + AC > AB 10 + 4 > sqrt(116); this correct AC + AB > sqrt(116) + 4 > 10 this correct we prrof that this traingle

OpenStudy (anonymous):

the second Question :- are this triangle is "Right triangle" ?

OpenStudy (anonymous):

if the traingle pythagorean theorem then the traingel is Right traingle

OpenStudy (anonymous):

\[(AB)^2 = (BC)^2 + (AC)^2\] 116 = 16 + 100 // then it's correct this is Right triangle